OpenStudy (anonymous):

one from linear programming : Minimise Z = −3x + 4y

OpenStudy (anonymous):

Subject to:

OpenStudy (anonymous):

\[\text{Minimize}[\{-3x+4y,x+2y\leq 8,3x+2y\leq 12,x\geq 0,y\geq 0\},\{x,y\}] \]\[\{-12,\{x\to 4,y\to 0\}\} \]where -12 is the minimum value of -3x+4y Was the preceding helpful?